### v1

# Does GPT Really Get It? A Hierarchical Scale to Quantify Human and AI’s Understanding of Algorithms

Abstract
As Large Language Models (LLMs) perform (and sometimes excel at) more and more complex cognitive tasks, a natural question is whether AI really understands. The study of understanding in LLMs is in its infancy, and the community has yet to incorporate well-trodden research in philosophy, psychology, and education. We initiate this, specifically focusing on understanding algorithms, and propose a hierarchy of levels of understanding. We use the hierarchy to design and conduct a study with human subjects (undergraduate and graduate students) as well as large language models (generations of GPT), revealing interesting similarities and differences. We expect that our rigorous criteria will be useful to keep track of AI’s progress in such cognitive domains.
1 Introduction
Since the release of GPT-4, mainstream users have begun to rely on Large Language Models (LLMs) for increasingly complex tasks. However, the degree to which it is safe, legal, and ethical to rely on LLMs has been under fierce debate. Across many studies, researchers have identified skills at which LLMs appear to fall short, such hallucinations, failure to plan, and lack of understanding (Rawte, Sheth, and Das 2023; Mahowald et al. 2024; Valmeekam et al. 2023). However, the literature notably lacks rigorous criteria to measure the progress toward solving these issues. A particular problem lies in claims surrounding understanding; AI understanding is frequently compared to human understanding, and it is folklore among AI researchers that the reasoning process of LLMs differ from humans. However, while the concept of understanding is widely discussed, it remains ill-defined.
In this paper, we propose an algorithmic definition of understanding an algorithm with the following properties: (a) it provides a scale by which to evaluate any agent’s understanding of an algorithm, (b) aligns with the standard usage of the term ‘understanding’ in philosophy and psychology, and (c) can evaluate LLM’s progress toward understanding algorithms.
Motivation - Why study algorithm understanding?
Large language models are increasingly trusted for coding assistance. Code generation tools such as GitHub Copilot (GitHub 2024) and Meta’s Code Llama (Roziere et al. 2023) are currently used in practice to improve developer productivity(Vaithilingam, Zhang, and Glassman 2022; Mozannar et al. 2024) and assist novice programmers in learning(Kazemitabaar et al. 2023; Becker et al. 2023). It is likely that the degree of AI involvement in software development will only grow as these tools improve. However, reliance on imperfect systems comes with risk. Tools such as Copilot are known to generate code that is subject to license(Becker et al. 2023) or contains security vulnerabilities(Pearce et al. 2022).
We argue that the question of whether LLMs demonstrate meaningful understanding of algorithms is relevant if we are relying on LLMs implement algorithms in production or teach them to novice programmers.
We additionally argue that algorithm understanding is a distinct concept from language understanding that deserves its own line of study. Those who argue that LLMs do not understand language draw a distinction between linguistic form and meaning(Bender and Koller 2020; Mitchell and Krakauer 2023; Pavlick 2023). When humans understand language, their understanding is informed by their communicative intent and the real-life properties of the objects described. Thus, a system trained only to replicate statistical correlations between words cannot understand language in the way that humans do. Algorithms, however, can be precisely represented using formal programming languages. One could argue that a computer can meaningfully observe an algorithm in full through code implementations and examples.
That is not to say that a human’s understanding of algorithms is not informed by real-world experience. For example, a student may understand shortest-path algorithms partially through their experience using navigational apps to get to work. Nonetheless, it is easier to work around this discrepancy between humans and LLMs when the object of study is formally defined.
Related Work
Cognitive Abilities of LLMs.
The past few years have seen an explosion of studies exploring the ability of LLMs to answer complex mathematical questions. Researchers have developed prompting strategies to enable multi-step reasoning (Wei et al. 2022; Fu et al. 2022). Still others fine-tune models to improve mathematical provlem-solving (Yu et al. 2023; Luo et al. 2023). The benchmarks for these methods typically include large datasets such as GSM8k (Cobbe et al. 2021) (grade school word problems) and MATH (Hendrycks et al. 2021) (math competition problems). These works focus on correct evaluation and do not address whether the language models understand mathematical reasoning.
Others have studied metacognitive skills in LLMs. Didolkar et al. (2024) investigate whether LLMs can assign skill labels to mathematical problems. Also related is Aher, Arriaga, and Kalai (2023) which proposes Turing experiments comparing humans and LLM simulations.
Understanding in LLMs
A parallel line of work investigates language understanding in LLMs. A key concept in the debate over language understanding is the difference between linguistic form and meaning (Bender and Koller 2020; Merrill et al. 2021). Bender and Koller (2020) argue that an AI trained only on linguistic form (i.e. text) cannot understand meaning. In an opinion piece, Pavlick (2023) counters this perspective, arguing that it is premature to draw conclusions on whether LLMs can model language understanding when the study of language models is itself in its infancy. There has been some effort to determine the extent to which LLMs represent linguistic meaning, primarily by studying word representations (Li, Nye, and Andreas 2021; Patel and Pavlick 2021). For a survey on linguistic competence in LLMs, see (Mahowald et al. 2024). Also see (Mitchell and Krakauer 2023) for a general survey on the debate over understanding.
Theories of Understanding
The debate over what constitutes understanding has a long history in philosophy and psychology. It is generally agreed that understanding is different from ‘mere’ knowledge, but the nature of that distinction is up for debate (Pritchard 2009; Baumberger, Beisbart, and Brun 2016; Páez 2019). Pritchard (2014) provides some examples of when the concepts of ‘knowing why’ and ‘understanding why’ may not overlap. Khalifa (2017) and Baumberger, Beisbart, and Brun (2016) are accessible surveys of this debate.
The philosophy of science also relates understanding and explanation, and the goal of explanation can be thought of as the production of understanding (Friedman 1974; Grimm 2010; Baumberger, Beisbart, and Brun 2016). Wilkenfeld, Plunkett, and Lombrozo (2016) argue the converse; they relate understanding to explanatory depth and claim that we attribute understanding in order to identify experts to consult. Woodward (2005) overviews what defines a causal explanation.
Also relevant to this work is the distinction between deep and surface-level learning from educational psychology (Marton and Säljö 1976; Beattie IV, Collins, and McInnes 1997). One influential categorization comes from Mayer (2002), who categorized student learning into cognitive processes and identified testable skills which arise with understanding. This categorization partially informed the levels we define below.
2 An Algorithmic Definition
We ask the question: how well does an agent understand an algorithm? Our goal is to create a definition of understanding that is testable and algorithmic. Therefore, we adopt a functional lens, meaning that we define understanding by what it allows the agent to do.
Preliminaries
In this work, we ask whether an agent $\mathcal{E}$ understands a computable function $f:\Omega\rightarrow\Sigma^{*}$. By computable, we say that there exists a Turing Machine which takes $x\in\Omega$ as input and halts with $f(x)$ on its tape, using a standard definition (Sipser 1996). Let $\mathcal{A}$ be an algorithm that computes $f$.
We assume that $\mathcal{E}$ (1) has a long-term memory system and (2) can perform computation, enabled by a working memory with finite capacity $M_{\mathcal{E}}$. We say that $\mathcal{E}$ knows a function $f$ or algorithm $\mathcal{A}$ if it has a representation $R_{f}/R_{\mathcal{A}}$ stored in its long-term memory. When $\mathcal{E}$ computes the function $f$ on an input $x\in\Omega$, it runs an internal algorithm $\mathcal{A}_{\mathcal{E}}$, which may or may not be the same as $\mathcal{A}$. The agent’s understanding of $f$ or $\mathcal{A}$ will be measured by its ability to manipulate this representation to produce answers to queries. This definition is based on the Understanding as Representation Manipulability system (URM) proposed by Wilkenfield (Wilkenfeld 2013). This system and the form of the representation is discussed further in the next section.
Internal Representations
When executed by a computer, an algorithm is represented as a string of bytes that are interpreted as a series of instructions. These instructions are stored in the computer’s memory, and in this way we can say that the computer has a representation of the algorithm.
Both humans and language models also have representations of the algorithm stored in their memory. For language models, this representation is collected from the thousands of examples, explanations, and code snippets that appear in its training data. The mechanism behind human memory is not understood as precisely. However, humans also learn via hearing explanations, collecting examples, and reinforcing their knowledge. This forms a representation encoded in the neural pathways of our brains (Durstewitz, Seamans, and Sejnowski 2000; György Buzsáki 2019)
We employ the framework of Understanding as Representation Manipulability (URM)(Wilkenfeld 2013). This theory posits that understanding arises from the ability to modify the internal representation of a concept in order to make effective inferences. In our example above, while a computer possesses a representation of the algorithm, it cannot independently modify it to answer questions. On the other hand, while a human may make more mathematical mistakes, they can answer interesting questions by ‘turning the problem around’ in their head.
Wilkenfeld declines to characterize the structure of the structure of the representations, other than to state that they are “computational structures with content that are susceptible to mental transformations”(Wilkenfeld 2013). Because of this, it is a general system that is applicable to many different types of agents.
It is interesting to think of a language model in this framework because it is clear what the representation and modification should be. The representation is a vector of weights, and a question related to the algorithm is also encoded as weights. The LLM makes inferences by performing mathematical operations on this vector (in essence, manipulating it).
Levels of Understanding
In this section, we define understanding as a spectrum by presenting a series of levels. Understanding at each level is loosely more difficult than the previous one, although they do not formally follow each other. Rather, they measure increasing levels of abstraction. To demonstrate the ideas, we also provide examples of questions which could be successfully answered by an agent that understands the Euclidean algorithm at each level.
For simplicity, we present these levels as deterministic; however, they can be defined with a failure probability dependent on the agent’s internal randomness and the required memory and time.
At the first level (denoted Level 1), the agent is capable of evaluating the algorithm on some ‘simple’ examples. At this level, the agent has some representation of the input-output mapping, whether or not it can formally express it.
Definition (Level 1: Execution).
$\mathcal{E}$ understands $\mathcal{A}$ at Level 1 if there exists parameters $M_{0},T_{0}$ such that the following holds: for any $x\in\Omega$ with $M(x)\leq M_{0}$ and $T(x)\leq T_{0}$, $\mathcal{A}_{\mathcal{E}}(x)=f(x)$.
Example: Compute GCD(24, 15).
At the next level, the agent can describe how it evaluates $f(x)$ in a language that it knows. Level 2 requires the agent to output the execution steps of the algorithm on $x$ as well as produce the correct answer.
Definition (Level 2: Step-By-Step Evaluation).
$\mathcal{E}$ understands $\mathcal{A}$ at Level 2 if, given an $x\in\Omega$ with $M(x)\leq M_{\mathcal{E}}$:
  • a list of instructions
  • a description in structured natural language
  • a flow chart or other unambiguous pictoral representation of the instructions
that are run when executing $\mathcal{A}$ on $x$.
Example: Compute GCD(462, 948) and show each step of the calculation.
The next level will take this one step further, requiring the agent to produce a set of instructions that can be followed to produce the right answer for any $x\in\Omega$.
Definition (Level 3: Representation).
$\mathcal{E}$ understands $\mathcal{A}$ at Level 3 if it understands at Levels 1 and 2, and it can produce one of the following:
  • a formal representation; e.g., code for $\mathcal{A}$ in a Turing-complete programming language it knows, a structured natural language description, an abstract syntax tree or Turing machine diagram.
  • an unambiguous description of the execution steps in natural language.
Example: Write a function in a programming language you know that can compute the GCD of any two integers.
The first three levels measure the ability of the agent to recall a procedure and execute a known set of instructions. We place these in the category of ‘shallow learning’; in Mayer’s taxonomy, they fall under the cognitive processes of recognizing, recalling, and executing (Mayer 2002). Note that all three levels could be achieved by a hard-coded script.
The next two levels target deep learning, and measure cognitive processes in the ‘Understanding’ and ‘Analyzing’ categories. We split the next levels into two subtrees, broadly corresponding to examples and explanation. This approach allows us to distinguish cognitive processes utilizing functional linguistic skills from those utilizing mathematical reasoning (Mahowald et al. 2024).
At Level 4, the agent can generate examples, convert between representations, and produce descriptions at different levels of detail.
Definition (Level 4a: Exemplification).
Given a property $P$ of the execution path of the algorithm, $\mathcal{E}$ can generate an $x\in\Omega$ which satisfies $P$ or report that none exists. Example: Give an integer $550<x<$ 55 that requires the greatest number of recursive steps to compute GCD$(55,x)$. Describe how you chose this number.
Definition (Level 4b: Summarization).
$\mathcal{E}$ can produce descriptions of $\mathcal{A}$ satisfying the following criteria.
  • When given a portion of the code for $\mathcal{A}$ in a formal language that it knows, the agent can summarize it in natural language. Similarly, it can produce correct, executable code when given a summary of a portion of the algorithm.
  • The agent can produce natural language summaries of the algorithm at distinct levels of abstraction. A more abstract summary replaces sections of the control flow with references to other algorithms $\mathcal{E}$ can describe.
Example: You are teaching a student who understands basic math operations but struggles with algebra and division with remainders. Explain how the Euclidean algorithm is used to find the greatest common divisor (GCD) of two given numbers, prioritizing intuition.
At Level 5, the agent can reason on perturbations of the algorithm.
Definition (Level 5a: Extrapolation).
Given an algorithm $\mathcal{A^{\prime}}$, the agent can determine whether $\mathcal{A}$ and $\mathcal{A^{\prime}}$ produce the same output on all $x\in\Omega$. If not, it can find a counterexample such that $\mathcal{A}^{\prime}(x)\neq f(x)$.
Example: Suppose GCD$(a,b)=x$ and GCD$(b,c)=y$. Does GCD$(a,b,c)$ = GCD$(x,y)$? Explain your reasoning.
Definition (Level 5b: Counterfactual Reasoning).
Given an algorithm $\mathcal{A^{\prime}}$, $\mathcal{E}$ can answer natural language queries about $\mathcal{A^{\prime}}$ phrased as counterfactual questions about $\mathcal{A}$.
Example: Consider the Fibonacci sequence defined by $F(0)=0$, $F(1)=1$, and $F(n)=F(n-1)+F(n-2)$ for $n\geq 2$. Why do consecutive Fibonacci numbers result in the maximum number of iterations for the Euclidean algorithm?
A counterfactual question asks ‘what if?’ and presents a modification to the basic assumptions of the setting. Counterfactual questions are considered in the philosophy of understanding to separate ‘knowing why’ from ‘understanding why’(Woodward 2005; Páez 2019). Páez (2019) gives the example of understanding climate change. A person might know that global warming is caused by greenhouse gases being released into the atmosphere, and could give a well-reasoned explanation of this phenomenon when asked. However, the same person may not be able to answer questions such as “what would happen to global temperatures if all human activity were to cease?” However, a climatologist, who likely understands why global warming occurs, could reason about this alternate scenario and provide an answer.
In terms of algorithms, we interpret this as the ability to answer questions about hypothetical modifications to the procedure, the input space, or the output space. In the example given, $\mathcal{A^{\prime}}$ restricts the input space to consecutive Fibonacci numbers. The task requires the agent to reason about changes in runtime under this restricted space. These levels of understanding and the partial ordering among them is summarized in Fig. 1.
Figure 1: A hierarchy of understanding.
3 Methods
We use two classical algorithms to test our scale of understanding: the Euclidean algorithm for computing the greatest common denominator of two integers, and the Ford-Fulkerson algorithm for computing the maximum flow between two nodes on a directed graph with capacity constraints (Ford and Fulkerson 1956). Both algorithms are widely taught in undergraduate computer science curricula. The Euclidean algorithm is relatively simple, and it is the first introduction to the concept of algorithm for many students. The Ford-Fulkerson algorithm, on the other hand, is more sophisticated and is typically introduced in a full-fledged algorithms course.
Human Survey
We conducted a survey on students of algorithms courses at a premier CS-teaching university. Each student was assigned either the Euclidean or Ford-Fulkerson Algorithm at random, and was asked to rate their own understanding of the algorithm on a six-point scale. Each survey consisted of five test questions to test their understanding. There were three versions of the survey for each algorithm, assigned at random. The questions are available in Appendix A.
The number of participants in the survey was $n=34$ (10 doctoral and 24 undergraduate). Students who reported that they did not understand the algorithm or completed less than half of the survey questions were removed the analysis. This left $n=23$ students (10 doctoral and 13 undergraduate). Of these students, ten had some teaching assistant experience in algorithms classes.
The questions are designed to map to increasing levels of understanding.
  1. 1.
    A trivial instance of the problem. If the student understands the input and output space, the problem can be answered without calculation.
  2. 2.
    An intermediate instance of the problem. This is answerable without a calculator for most undergraduates, but requires the student to run some internal algorithm to compute.
  3. 3.
    A coding problem. This problem requires the student to translate a part of the algorithm to code.
  4. 4.
    Either an example problem or an explanation problem. This question asks the student to provide an example to illustrate a concept or explain the algorithm to a specified audience.
  5. 5.
    A counter-example or extension problem. This problem asks the student to reason about modifications to the algorithm.
LLM Experiments
Several LLM models were presented with the same surveys given to the human participants; each survey was started in a fresh chat session, and the five questions were presented in the same order (within the survey, previous questions and responses were included in the chat history). We also included a system prompt to prime the LLM with the algorithm in question and encourage conciseness in the responses.
To ensure the statistical stability of the results, the LLM was also queried using randomized versions of the survey. For evaluation questions, the values in the questions were assigned uniformly at random within a given range. For the flow questions, the graph structure was also varied slightly. For the code questions, we took an example code implementation of the algorithm, randomly masked a line or group of lines, and asked the LLM to fill in the missing part. Lastly, for the example, explanation, and extension questions, we automatically generated variants of the question using GPT-4o (in a separate chat session). These questions were manually inspected to ensure they matched our criteria, and questions which were not useful were discarded.
Three versions of GPT were tested: GPT 3.5 Turbo, GPT 4, and GPT-4o.
Evaluation
Each question is rated on a scale from zero to two. With the exception of the explanation questions, the scores have the following interpretations: (0) incorrect; (1) partially correct, surface level; (2) completely correct, thorough.
Evaluating Explanations
The quality of explanations and summaries can be subjective; however, they offer a deep insight into the subject’s understanding of the material. We evaluate the correctness of explanations on three axes.
  1. 1.
    First, correctness; the explanation is accurate and includes the key ideas of the algorithm.
  2. 2.
    Second, audience adaption; the explanation is tuned to the knowledge and needs of the audience.
  3. 3.
    Third, intuitiveness; the explanation conveys intuition - for example via contrast, example, or analogy - and uses clear language.
Note that summaries and explanations are by definition selective, and are not necessarily complete (Mittelstadt, Russell, and Wachter 2019). In fact, the ability to identify the key ideas are part of what differentiates explanation (Level 4) from the production of instructions (Level 3). The explanation is awarded 2/3 of a point for each bullet, for a maximum score of 2 per question.
Hypotheses
1. The understanding hierarchy (Figure 1) captures depth of understanding.
Therefore, any student who answers a question at a later level is likely to be able to answer a question from an earlier level. As a consequence, the fraction of correct responses should be monotonically non-increasing with the level, and the conditional probability of a correct response for a question should be higher if the previous question was answered correctly.
2. The landscape of understanding for GPT will be different from that of humans.
Since explanations of the algorithms and their properties appear frequently in the training data, we expect GPT-4 to be successful at questions related to explanation and summary. For the same reason (exposure to large corpora of code in training data), we expect it to do well at coding questions. However, we expect it to struggle with execution, including Questions 1 and 2, which require calculation. We expect this to be true for to a greater extent for GPT-3.5, which appears to have weaker arithmetic abilities.
3. The performance of GPT-4 can be improved via techniques such as chain-of-thought prompting and the inclusion of example solutions. These are recommended practices (OpenAI 2024).
4 Results
Out of all the responses collected, 85% of students indicated that they understood the algorithm, with most students reporting that they “know the algorithm and have a fair understanding of it”. For further analysis, we only consider responses from students who stated that they understood the algorithm. Across all students, the accuracy on the questions was highest for Question 1, and decreased uniformly through Question 5 (Figure 2).
Among versions of GPT, the best performing version was GPT-4 (Figure 4). The response score of GPT-4 was close to that of graduate students, as shown in Figure 2 (although this statement has some caveats, discussed further under Qualitative Comparison). LLMs on average out-performed the undergraduate students, except on questions 1 and 2 (evaluation). We take a closer look at the score distribution in Figure 5. The LLM performance exhibited a similar accuracy curve to students, with the exception of question 3 - all three versions of GPT tested were able to reliably answer the coding question regardless of whether the previous problem was evaluated correctly. This trend is exhibited in Figure 6.
Figure 2: The average score per question across undergraduate and doctoral students who self-reported that they understood the algorithm. Number of records is $n=13$ and $n=10$ respectively. Also shown is the average score by GPT-4 across 57 randomized versions of the surveys.
Figure 3: The average score per question and algorithm across undergraduate and graduate students who self-reported that they understood the algorithm.
Figure 4: The average score between three versions of GPT, across 27 randomized surveys on Max Flow and 30 on GCD.
Figure 5: The distribution of scores per question for GPT-4.
Figure 6: Difference in score between questions at consecutive levels, for both students (orange) and versions of GPT (blue). The circles are the means and the bars are the standard deviations.
Alignment with Levels.
In Figure 7, we map the results back to the levels shown in Figure 1. We computed the average score across questions in each category for each algorithm. Both GPT 4 and GPT 4o demonstrated an increase in quality in execution (Levels 1 and 2), explanation (Level 4b), and extension (Level 5b). On the Ford-Fulkerson algorithm, it’s ability to produce examples improved only marginally, and no version of GPT was able to successfully answer the difficult counterexample questions.
(a) GPT 3.5 Turbo
(b) GPT 4
(c) GPT 4o
(d) Undergraduate ($n=5$)
(e) Doctoral ($n=5$)
(f) GPT 3.5 Turbo
(g) GPT 4
(h) GPT 4o
(i) Undergraduate ($n=8$)
(j) Doctoral ($n=5$)
Figure 7: Average score at each level on the Ford Fulkerson Algorithm (Top, a-e) and Euclidean Algorithm (Bottom, f-j) for each GPT version. Darker colors indicate higher average scores for that question category.
Prompting with examples.
[…]
Figure 8: Percent accuracy on 100 randomized max flow problems per GPT version. Shows both base accuracy (blue) and accuracy when a correct example was given in the chat history (orange).
We find that this strategy generally causes the response to mimic the structure of the example response. As shown in Figure 8, including an example response improves accuracy for GPT 3.5, but has little-to-no effect on GPT 4o, and marginally decreases the accuracy for GPT 4. One possible explanation for this phenomenon is that GPT 4 naturally responds to the prompt with effective chain of thought reasoning; therefore, instructing it to reason in a specific format does not improve its reasoning abilities, and may in fact interfere with them.
Qualitative Comparison.
This section will highlight patterns in responses that cannot be seen in the aggregate results.
Generating Examples. As shown in Figure 7, all versions of GPT struggled to produce an example of a graph satisfying a prescribed property. This could be attributed to GPT’s known difficulties with mathematical calculation - after all, in order to know if the execution of the algorithm satisfies a property, it must first execute the algorithm, which GPT generally cannot do without chain-of-thought prompting(Wei et al. 2022). However, our evidence suggests a deeper issue with categorizing inputs.
We asked the following question: “Give an example of a graph where the Ford-Fulkerson algorithm computes exactly six augmenting flows before terminating. Write the example as a list of edges and capacities.” This question is deliberately ambiguous, as the number of augmenting flows depends on the method used to find paths. This ambiguity, along with the relatively large number of paths, should steer the respondent away from a ‘trial and error’ method. The intended answer is a graph with a source s, a sink t, and six intermediate vertices connected to both, leading to six parallel paths from s to t.
In the human survey, out of nine students, six gave this exact answer (the other three did not attempt the question). It is reasonable to expect that these students did not mentally execute the Ford-Fulkerson algorithm in order to verify their answer. Instead, they likely had some knowledge of the concept of parallel paths, and were able to leverage this knowledge to retrieve an example.
The graphs given by GPT did not follow any discernible pattern. Over all trials, none required six augmenting paths (all examples reviewed were too small, admitting at most 5 augmenting paths regardless of the path-finding algorithm). We argue that this suggests that GPT lacks an ability to manipulate its representation of the input space to produce useful shortcuts.
The Fibonacci Sequence. We asked two variant questions exploring the relationship between the Fibonacci sequence and the Euclidean algorithm. One question asked the respondent to explain why consecutive Fibonacci numbers are considered worst-case inputs for the Euclidean algorithm. A correct answer would point out that since $F(n+1)\text{mod}F(n)=F(n-1)$, the input size reduces by approximately $\phi=1.618\dots$ at each step, leading to the worst-case run time of $O(\log(a))$. The second question asked the respondent to find an integer $x$ such that computing $\text{gcd}(55,x)$ takes the maximum number of steps, and to explain their choice. This question did not include the word Fibonacci, and the two variants were never asked to the same student so as not to bias the result.
Out of the ten students asked the second question, four provided a correct response. Although the word Fibonacci was never stated, a few students seemed to intuit the motivation behind the question. One doctoral student responded:
It should be approximately 1/phi (golden ratio) times 55 in general I think. Since 55 is a Fibonacci number it is easy to guess that the longest sequence should be obtained by 34, the preceding Fibonacci number, since this results in the Euclidean algorithm going through the Fibonacci sequence down to 1.
Another stated,
55 is a fibonacci number so I believe the answer is x=34 although I am not sure. That would require you to take one step in each direction each time.
This question also highlighted a jump in performance from GPT 3.5 to GPT 4. GPT 3.5 got an average score of 0.5/2 on the question, while GPT 4 and 4o averaged 1.9 and 2.0 respectively. However, GPT 3.5 also seemed to ‘intuit’ that the question involved the Fibonacci sequence. In five out of twelve responses, GPT 3.5 responded with the correct answer $x=34$, but with an incomplete or incorrect explanation (correct answers with incomplete explanations were scored as 1.33/2, while correct answers with nonsensical explanations were scored as 0.67/2). One response read:
The number 34 in the interval 0 $<$ x $<$ 55 leads to the maximal number of recursive steps. This is because 55 and 34 are consecutive Fibonacci numbers, which tend to lead to the largest number of recursive steps in the Euclidean algorithm due to their relatively prime nature.
While it is true that consecutive Fibonacci numbers are relatively prime, this does not explain why it requires more recursive steps than other pairs of numbers (e.g., while 54 and 55 are relatively prime, computing GCD(54, 55) only takes two recursive steps). Other explanations given by GPT were completely incorrect:
The integer $x$ that results in the longest sequence of recursive calls when calculating GCD($55,x$) using the Euclidean algorithm is $x=34$.
Explanation: […]. In this case, since $34$ is the largest integer less than $55$ that has a common factor with $55$ (GCD(55, 34) = 1), it results in the longest sequence of recursive calls compared to other integers. This is because the smaller the common factor, the more iterations are needed to reach a remainder of 00.
Responses such as this support the theory that the production of a correct answer is not direct evidence that LLMs have understanding.
Trouble with Ropes. For both the algorithms, one question asked the respondent to explain the algorithm to a student struggling with certain key concepts. One tactic which emerged was to explain the algorithm using a physical analogy. Thirteen out of the twenty-eight LLM responses asked the student to imagine the Ford-Fulkerson algorithm using water flowing through a series of pipes, and ten out of twenty-five LLM responses described the Euclidean algorithm as a way to find the smallest piece of rope that equally divides two ropes or sticks of different lengths. In contrast, one out of five student responses described FF using a series of pipes, and none of the six student responses used a physical analogy to explain the Euclidean algorithm.
We found three out of ten responses from LLMs using the rope/stick analogy for the Euclidean algorithm included explanations which did not make physical sense. For example, one response from GPT 3.5 described the algorithm as follows:
Take the two rope lengths (let’s call them Rope A and Rope B), and see how many times the shorter rope (Ride (sic) A) can wrap around the length of the longer rope (Rope B) without hanging over.
It is obvious to a human reader that one cannot wrap the shorter rope around the longer. This matches findings that LLMs lack functional understanding of properties of real objects (Mahowald et al. 2024).
Hallucinations. Within correct responses, we found that large language models included nonsequitors and inaccurate statements. As mentioned above, one extension question asked why consecutive Fibonacci numbers are considered ‘worst case’ inputs for the Euclidean algorithm. GPT-4 got a score of 2/2 on this question 92% of the time. However, within a correct answer, it occasionally included incorrect statements. For example, in one response it included the following line (emphasis ours):
1. **Structure of Fibonacci numbers**: Each Fibonacci number is the sum of the two preceding ones, with the sequence beginning as $F(0)=0$, $F(1)=1$, $F(2)=1$, $F(3)=2$, and so forth. This means each number in the sequence is relatively close to the sum of the two preceding numbers.
In another response, it stated that $\phi$ was “one of the most irrational numbers”. This highlights some need for caution when using GPT-4 as a teaching implement.
5 Discussion
We have presented a hierarchical scale for measuring understanding of algorithms in both humans and AI. In this section, we discuss implications, interpretations amd limitations of our approach and point to future directions.
Our results show a significant improvement between the GPT 3.5 and GPT 4 generations in the skills of evaluation, explanation, and counterfactual reasoning for the two tested algorithms.
All versions of GPT performed near-perfectly on the skill of representation. This is in line with other findings showing the competence of GPT in code generation (Vaithilingam, Zhang, and Glassman 2022; Savelka et al. 2023). On the flip side, this suggests that if an LLM is able to generate code, it may not necessarily be able to perform other reasoning tasks such as evaluation and explanation.
This result begs the question: is GPT actually reasoning, or can the responses be explained by more superficial statistical correlations? If many similar questions and answers can be found in its training data, then GPT may be able to produce correct answers by leveraging statistical correlations between the input and the correct response. Can this really be called understanding? To this we make two points. First, the evaluation questions are instantiated with random values, and the exact questions are almost certainly novel. We argue that this suggests that GPT must be making some nontrivial transformation to produce correct answers.
Second, we note that in order to determine whether GPT is reasoning, the concept of reasoning itself needs to be precisely defined. What our study indicates is that GPT-4 has a sophisticated internal representation, e.g., its representation of the Euclidean algorithm includes its relationship to Linear Diophantine equations and mathematical properties of GCD. It is able to retrieve these properties under a variety of contexts and framings. In a similar vein, it is likely that a student answering these questions has also been exposed to these properties in their studies. When evaluating the usefulness of a representation, we do not necessarily need to account for how the representation was created.
Limitations.
Our results do not imply that GPT-4 has, on average, a ‘better’ understanding of the tested algorithms than undergraduates. The student study was included as a basis of comparison, and to demonstrate that the hierarchy we present can also apply to humans. However, we work with a limited population size, and the difference is confounded by other factors such as subject fatigue. Further research is needed to compare the quality of GPT and human responses to questions pertaining to algorithms. We hope that our scale can be applied to other algorithmic and similarly precise realms of understanding.
Acknowledgements
The authors are grateful to Rosa Arriaga, Adam Kalai and Sashank Varma for helpful discussions. This work was funded in part by NSF Award CCF-2106444 and a Simons Investigator award.
Appendix A Survey Questions
This section contains the text of the survey given to human subjects. Each respondent was asked some preliminary questions about their course level and experience with algorithms. They were assigned either the Ford-Fulkerson algorithm or the Euclidean algorithm at random and asked to self-report their understanding (Figure 9). Some students who reported that they did not understand the Ford-Fulkerson algorithm were then given the survey for the Euclidean algorithm.
Figure 9: The six point-scale shown to students to self-report their understanding.
The student was then randomly assigned one of three survey versions, shown below.
Euclidean Algorithm
Version 1
  1. 1.
    Compute GCD(24, 15), and show each step of the algorithm.
  2. 2.
    Compute GCD(462, 946), and show each step of the algorithm.
  3. 3.
    Using a programming language you are familiar with, write code for a function gcd(a,b), which computes the greatest common denominator of two integers a and b using the Euclidean algorithm.
  4. 4.
    You are instructing a student in an algebra course. The student is familiar with basic mathematical operations such as addition, subtraction, multiplication and division. However, she struggles with algebraic equations and division with remainders. Explain the Euclidean algorithm to her, prioritizing conveying intuition.
  5. 5.
    Suppose gcd(a,b) = x and gcd(b,c) = y. Does gcd(a,b,c) = gcd(x,y)? Explain your reasoning.
Version 2
  1. 1.
    Compute GCD(24, 15), and show each step of the algorithm.
  2. 2.
    Compute GCD(4088, 1241), and show each step of the algorithm.
  3. 3.
    Using a programming language you are familiar with, write code for a function gcd(a,b), which computes the greatest common denominator of two integers a and b using the Euclidean algorithm
  4. 4.
    Consider computing GCD(55, x) for an input x. Give an integer $550<x<$ 55 that requires the greatest number of recursive steps to compute GCD(55, x). Describe how you chose this number.
  5. 5.
    Determine whether the following statement is true. If not, provide a counterexample. For any two positive, nonzero integers a,b, the equation sa + tb = GCD(a,b) has exactly one solution where s and t are integers.
Version 3
  1. 1.
    Compute GCD(24, 15), and show each step of the algorithm.
  2. 2.
    Compute GCD(1008, 468), and show each step of the algorithm.
  3. 3.
    Using a programming language you are familiar with, write code for a function gcd(a,b), which computes the greatest common denominator of two integers a and b using the Euclidean algorithm.
  4. 4.
    You are speaking with a mathematics student who understands modular arithmetic. Explain the proof that the Euclidean algorithm finds the greatest common denominator of two numbers.
  5. 5.
    The Fibonacci numbers are a sequence F(n) where F(0) = 0, F(1) = 1, and F(n) = F(n-1) + F(n-2) for any $n\geq 2$. Consecutive Fibonacci numbers can be thought of as ‘worst case’ inputs for the Euclidean algorithm. Can you explain why?
Ford-Fulkerson Algorithm
Version 1
  1. 1.
    Compute the maximum flow between A and B. List each augmenting path and the flow along the path at each step.
  2. 2.
    Compute the maximum flow between A and D. List each augmenting path and the flow along the path at each step.
  3. 3.
    The following is an implementation of breadth-first search in Python with one line missing. Write a condition to replace the highlighted text. If you are not familiar with Python syntax, you may use your best guess.
  4. 4.
    Give an example of a graph where the Ford-Fulkerson algorithm computes exactly six augmenting flows before terminating. Write the example as a list of edges and capacities.
  5. 5.
    Consider an implementation of Ford-Fulkerson which uses a breadth-first search to find the augmenting path. Give an example which illustrates why the algorithm needs to track residual capacity of reverse edges.
Version 2
  1. 1.
    Compute the maximum flow between A and B. List each augmenting path and the flow along the path at each step.
  2. 2.
    Compute the maximum flow between A and D. List each augmenting path and the flow along the path at each step.
  3. 3.
    The following is an implementation of Ford-Fulkerson in Python with two lines missing. Write code to replace the highlighted text. If you are not familiar with Python syntax, you may use your best guess.
  4. 4.
    You are a civil engineer at a city planning commission. Your boss has asked you whether it is possible to calculate the maximum volume of traffic that can be routed between two destinations in the city. Describe the Ford-Fulkerson algorithm, but since he is a busy man, do not bore him with the details.
  5. 5.
    Suppose you are an event planner, and you need to determine the maximum amount of traffic that can be routed from the airport to one of two event venues. In other words, you want to compute the maximum flow between a source s and two sinks t1 and t2. How would you implement this?
Version 3
  1. 1.
    Compute the maximum flow between A and B. List each augmenting path and the flow along the path at each step.
  2. 2.
    Compute the maximum flow between A and D. List each augmenting path and the flow along the path at each step.
  3. 3.
    The following is an implementation of Ford-Fulkerson in Python with three lines missing. Write code to replace the highlighted text. If you are not familiar with Python syntax, you may use your best guess.
  4. 4.
    You are assisting a student in an algorithms course. He is struggling to understand the Ford-Fulkerson algorithm for computing MAXFLOW. Describe it to him, prioritizing conveying intuition.
  5. 5.
    An s-t cut is a set of edges which, when removed, divide vertex s and vertex t into separate components. For example, the dotted edges in the graph below are an A-F cut with capacity 24.
    It turns out that the maximum s-t flow on a graph with capacities is equal to the minimum capacity of any s-t cut. Describe your intuition for why this might be true.
Appendix B Further details on LLM experiments
In each experiment on GPT, the query included a system prompt to encourage the model to produce only relevant information. The system prompts used in the experiments are as follows:
‘You will answer a series of questions related to computing the maximum flow on a directed graph. You provide concise responses and do not include detail or explanations unless explicitly requested by the user.’
‘You will answer a series of questions related to the Euclidean algorithm for computing the greatest common denominator (gcd) of two integers. You provide concise responses and do not include detail or explanations unless explicitly requested by the user.’
GPT was queried via the OpenAI Chat Completions API using the default parameters (e.g. temperature=1). Specific versions queried are ‘gpt-3.5-turbo-0125’,‘gpt-4-turbo-2024-04-09’, and ‘gpt-4o-2024-05-13’.
We conducted two experiments with GPT. In the first, GPT was asked five randomized versions of the questions in Section A (a quiz), with the images being replaced with text descriptions of the graphs. For each quiz, the questions corresponding to each level of understanding were asked in order, and the previous questions and responses were included in the API query. This quiz was repeated thirty times for each version. Three quizzes were discarded due to the responses to the randomized questions being incomparable with the others. All trials were hand-graded on a scale from zero to two using the scale described in Section 3.
In the second experiment, GPT was queried with only the simple and intermediate evaluation questions for Max Flow (with varying graph structure and random weights). In one trial, GPT was first asked the simple question followed by the intermediate (with its response to the simple question passed to the API in the chat history). In the second trial, a sample response to an intermediate evaluation was passed to the chat history. The responses were checked by hand for correctness, and the rate of correctness was recorded.